One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers. No book. Sorry! Thankyou!!!
Answer by funmath(2933)

Smaller number:x
Larger number:x+6
The sum (+) of the smaller number (x) and three times (3*) the larger number (x+6) is (=) 34
x+3(x+6)=34
x+3x+18=34
4x+18=34
4x+18-18=34-18
4x=16
4x/4=16/4
x=4
Smaller number=x=4
Larger number=x+6=4+6=10
---------------------------
Check:
The sum (+) of the smaller number (x) and three times (3*) the larger number (x+6) is (=) 34
4+3(10)=34
4+30=34